# Neural Diversity Regularizes Hallucinations in Small Language Models

### v1

@@ 6 Related Work
Hallucination in Language Models. Hallucinations represent a fundamental challenge in language model deployment (Huang et al., 2024; Tonmoy et al., 2024), proven mathematically inevitable in computable models (Xu et al., 2024; Kalai & Vempala, 2024), with particular severity in smaller models (Lin et al., 2021; Li et al., 2023a). Mechanistic investigations reveal hallucinations arise from internal representation failures (Yu et al., 2024; Ferrando et al., 2025). Current mitigation strategies—retrieval augmentation (Niu et al., 2024), specialized decoding (Li et al., 2023b; Manakul et al., 2023; Wei et al., 2024), and constitutional training (Bai et al., 2022)—address only specific aspects. In other contexts, MoE-LLaVA improves by 1.1% on POPE but requires $8\times$ parameters (Zhang et al., 2024). Our work achieves hallucination reduction through neural diversity within a single architecture at fixed parameter budgets.
Parallel Architectures and Scaling Laws. Parallel scaling offers a third axis beyond parameters and data (Chen et al., 2025; Kaplan et al., 2020). ParScale achieves $O(\log P)$ gains with 22× less memory and 6× less latency (Chen et al., 2025), while inference-optimal scaling laws favor smaller models with more data (Sardana & Frankle, 2024) and MoE achieves 1000× capacity through conditional computation (Shazeer et al., 2017). However, these approaches target accuracy rather than reliability. We show parallel scaling reduces hallucinations through explicit neural diversity.
[…]
Redundancy Reduction and Self-Supervised Learning. Self-supervised methods provide differentiable diversity mechanisms. Barlow Twins prevents collapse through dimension decorrelation (Zbontar et al., 2021), VICReg decomposes into variance, invariance, and covariance terms (Bardes et al., 2022), and understanding dimensional collapse (Jing et al., 2022) reveals why explicit decorrelation is necessary. We adapt these principles for hallucination reduction beyond visual learning.
Parameter-Efficient Fine-Tuning. PEFT methods enable distinct streams under fixed budgets. LoRA reduces parameters by 10,000× with multiple adapters inducing specialization (Hu et al., 2022; Wang et al., 2023), prefix-tuning uses 0.1% of parameters (Li & Liang, 2021), and BitFit shows minimal differences enable specialization when regularized (Ben Zaken et al., 2022). BatchEnsemble and LoRA-Ensemble achieve diversity through clever parameterization (Wen et al., 2020; Mühlematter et al., 2024), providing our technical foundation.
Inference-Time Scaling and Aggregation. Self-consistency improves reasoning by 17.9% through diverse sampling (Wang et al., 2022), though resampling has limitations with imperfect verifiers (Stroebl et al., 2024). Confidence-based weighting reduces required paths by 40% (Taubenfeld et al., 2025), while contrastive decoding and classifier-free guidance achieve improvements equivalent to parameter doubling (Li et al., 2023b; Sanchez et al., 2023). Unlike post-hoc aggregation requiring multiple forward passes, our training-time parallelism learns coordinated streams.
Theoretical Foundations. PAC-Bayesian bounds show minimax-optimal rates connecting diversity to generalization (Steffen et al., 2024; Biggs & Guedj, 2022; Ortega et al., 2022), while concentration inequalities demonstrate that reducing predictor correlation tightens tail bounds on error events (Alquier, 2024). This supports our approach of explicitly regularizing cross-stream correlation for factual reliability.
@@ 7 Discussion
Neural diversity establishes a principled framework bridging portfolio theory and deep learning to develop models that are architecturally robust to noise. Our work provides both theoretical grounding — proving hallucination probability is bounded by cross-stream correlation — and empirical validation through ND-LoRA’s 25.6% hallucination reduction at negligible computational cost. The framework reveals why existing scaling approaches target accuracy over reliability and how explicit decorrelation through neural diversity provides a mechanism to reduce hallucinations at fixed parameter budgets, introducing a third scaling axis orthogonal to parameters and data. Unlike traditional ensemble methods (Lakshminarayanan et al., 2017), which require multiple separate models, our approach achieves diversity benefits within a single architecture.
At the same time, our demonstration has limitations. Computational constraints restricted experiments to 0.5B models; larger-scale production-ready implementations are part of planned future work. Benchmark scope focuses on English hallucination and knowledge tasks, with some tasks showing no benefit from increased neural diversity, highlighting task-dependent efficacy that requires further investigation. Finally, while our theoretical framework predicts non-monotonicity, it doesn’t provide a satisfying explanation of why different tasks have different optimality points.
As language models become critical infrastructure, reliability techniques that avoid massive resource investment become essential. While hallucinations may be mathematically inevitable (Kalai & Vempala, 2024), neural diversity demonstrates they can be systematically reduced through architectural design. Our results show reliable AI may emerge from thoughtfully designed architectures rather than brute-force scaling alone, opening new avenues for efficient, trustworthy language models.
@@ Acknowledgments
We thank the research community at South Park Commons for valuable discussions and feedback throughout this project. We are grateful to Augustine Mavor-Parker, Bhav Ashok, Daniel Morillo, David Sontag, Iman Modarressi, Jaclyn Lunger, John Bohannon, Nelson Ray, Patrick Bozeman and Suhith Rajesh for their insightful comments on earlier drafts. This work was supported by South Park Commons and Modal, who generously provided compute resources for our experiments.
@@ A.1.1 Proof of Lemma 1
By bilinearity of covariance,
$\mathrm{Var}(M)$ $=\mathrm{Var}\!\Big(\frac{1}{P}\sum_{i=1}^{P}m_{i}\Big)=\frac{1}{P^{2}}\Bigg(\sum_{i=1}^{P}\mathrm{Var}(m_{i})\;+\;2\sum_{1\leq i<j\leq P}\mathrm{Cov}(m_{i},m_{j})\Bigg)$
$=\frac{1}{P^{2}}\Big(P\,\sigma^{2}\;+\;2\cdot\tbinom{P}{2}\,\rho\,\sigma^{2}\Big)=\frac{1}{P^{2}}\Big(P\,\sigma^{2}+P(P-1)\,\rho\,\sigma^{2}\Big)$
$=\sigma^{2}\!\left(\frac{1}{P}+\frac{P-1}{P}\,\rho\right)=\sigma^{2}\!\left(\frac{1-\rho}{P}+\rho\right).$
A.1.2 Proof of Lemma 2
@@
Since $m_{i}=v_{i}^{\top}\tilde{z}_{i}$ and $m_{j}=v_{j}^{\top}\tilde{z}_{j}$,
$\mathrm{Cov}(m_{i},m_{j})=\mathbb{E}\big[(v_{i}^{\top}\tilde{z}_{i})(v_{j}^{\top}\tilde{z}_{j})\big]=v_{i}^{\top}\mathbb{E}[\tilde{z}_{i}\tilde{z}_{j}^{\top}]v_{j}=v_{i}^{\top}C_{ij}v_{j}.$ (14)
By the definition of the spectral norm,
$|\mathrm{Cov}(m_{i},m_{j})|=|v_{i}^{\top}C_{ij}v_{j}|\leq\left\lVert v_{i}\right\rVert\,\left\lVert C_{ij}\right\rVert_{2}\,\left\lVert v_{j}\right\rVert.$ (15)
Divide both sides by $\sigma_{i}\sigma_{j}$ to obtain
$|\rho_{ij}|\leq\frac{\left\lVert v_{i}\right\rVert\,\left\lVert C_{ij}\right\rVert_{2}\,\left\lVert v_{j}\right\rVert}{\sigma_{i}\sigma_{j}}=\kappa_{ij}\,\left\lVert C_{ij}\right\rVert_{2}.$ (16)
Since the right-hand side is nonnegative, this yields the stated upper bound on $\rho_{ij}$. ∎
@@ A.1.3 Proof of Theorem 1
Let $\rho_{ij}=\mathrm{Corr}(m_{i},m_{j})$ and $\bar{\rho}\triangleq\frac{2}{P(P-1)}\sum_{i<j}\rho_{ij}$. With common per-stream variance $\sigma^{2}$,
$\mathrm{Var}(M)$ $=\mathrm{Var}\!\Big(\frac{1}{P}\sum_{i=1}^{P}m_{i}\Big)=\frac{1}{P^{2}}\Big(\sum_{i}\mathrm{Var}(m_{i})\;+\;2\sum_{i<j}\mathrm{Cov}(m_{i},m_{j})\Big)$
$=\frac{1}{P^{2}}\Big(P\sigma^{2}+2\textstyle\binom{P}{2}\sigma^{2}\,\bar{\rho}\Big)=\sigma^{2}\!\left(\frac{1}{P}+\frac{P-1}{P}\,\bar{\rho}\right)=\sigma^{2}\!\left(\frac{1-\bar{\rho}}{P}+\bar{\rho}\right).$
Next, by the linear feature model $m_{i}=v_{i}^{\top}\tilde{z}_{i}$ and the definition of $C_{ij}=\mathbb{E}[\tilde{z}_{i}\tilde{z}_{j}^{\top}]$,
$\mathrm{Cov}(m_{i},m_{j})=v_{i}^{\top}C_{ij}v_{j},\qquad|\mathrm{Cov}(m_{i},m_{j})|\leq\|v_{i}\|\,\|C_{ij}\|_{2}\,\|v_{j}\|.$ (17)
Dividing by $\sigma_{i}\sigma_{j}$ gives
$\rho_{ij}\;=\;\frac{\mathrm{Cov}(m_{i},m_{j})}{\sigma_{i}\sigma_{j}}\;\leq\;\frac{\|v_{i}\|\,\|v_{j}\|}{\sigma_{i}\sigma_{j}}\,\|C_{ij}\|_{2}\;=\;\kappa_{ij}\,\|C_{ij}\|_{2}.$ (18)
Averaging over unordered pairs,
$\bar{\rho}=\frac{2}{P(P-1)}\sum_{i<j}\rho_{ij}\;\leq\;\frac{2}{P(P-1)}\sum_{i<j}\kappa_{ij}\,\|C_{ij}\|_{2}=\bar{\kappa}\cdot\frac{2}{P(P-1)}\sum_{i<j}\|C_{ij}\|_{2}=\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}.$
Because $\mathrm{Var}(M)$ is increasing in $\bar{\rho}$ for $P>1$,
$\mathrm{Var}(M)\;\leq\;\sigma^{2}\!\left(\frac{1-\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}}{P}+\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}\right).$ (19)
Applying Cantelli’s inequality to $M$ with $\mathbb{E}[M]=\mu>0$ and $t=\mu$,
$\mathbb{P}(M\leq 0)=\mathbb{P}(M-\mu\leq-\mu)\leq\frac{\mathrm{Var}(M)}{\mathrm{Var}(M)+\mu^{2}}\leq\frac{\sigma^{2}\left(\tfrac{1-\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}}{P}+\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}\right)}{\sigma^{2}\left(\tfrac{1-\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}}{P}+\bar{\kappa}\,\mathcal{D}_{\mathrm{spec}}\right)+\mu^{2}}.$ (20)
Finally, if assumptions (ii)–(iii) hold only on a “good” event $G$ with $\mathbb{P}(G)\geq 1-h_{0}$, then
$\mathbb{P}(M\leq 0)=\mathbb{P}(M\leq 0\mid G)\,\mathbb{P}(G)+\mathbb{P}(M\leq 0\mid G^{\mathrm{c}})\,\mathbb{P}(G^{\mathrm{c}})\leq\mathbb{P}(M\leq 0\mid G)+h_{0},$ (21)
and the displayed fraction bounds $\mathbb{P}(M\leq 0\mid G)$, yielding the stated result. ∎
A.1.4 Proof of Theorem 2
Proof.
Let $g(P)\triangleq\tfrac{1-\rho(P)}{P}+\rho(P)$ so that $\mathcal{B}(P)$ is an increasing function of $g(P)$ (monotonicity in the numerator). Compute
$g^{\prime}(P)=\frac{d}{dP}\!\left(\frac{1-\rho(P)}{P}\right)+\rho^{\prime}(P)=\Big(-\frac{\rho^{\prime}(P)}{P}-\frac{1-\rho(P)}{P^{2}}\Big)+\rho^{\prime}(P)=\rho^{\prime}(P)\Big(1-\frac{1}{P}\Big)-\frac{1-\rho(P)}{P^{2}}.$ (22)
As $P\downarrow 1$, we have $1-\tfrac{1}{P}\to 0$ while $1-\rho(P)\to 1-\rho_{0}>0$, hence $g^{\prime}(P)\to-(1-\rho_{0})<0$. For large $P$, the negative term is $O(P^{-2})$, while $\rho^{\prime}(P)=\beta\gamma(P-1)^{\gamma-1}>0$; thus for sufficiently large $P$, $g^{\prime}(P)>0$. By continuity, there exists $P_{\star}>1$ with $g^{\prime}(P_{\star})=0$, implying that $g$ (and therefore $\mathcal{B}$) decreases for $P<P_{\star}$ and increases for $P>P_{\star}$.
If $\gamma\geq 1$, then $\rho^{\prime\prime}(P)=\beta\gamma(\gamma-1)(P-1)^{\gamma-2}\geq 0$ and
$g^{\prime\prime}(P)=\rho^{\prime\prime}(P)\Big(1-\frac{1}{P}\Big)+\frac{2\rho^{\prime}(P)}{P^{2}}+\frac{2(1-\rho(P))}{P^{3}}\;\geq\;0\quad\text{for }P>1.$ (23)
Hence $g$ is convex on $(1,\infty)$ and has a unique minimizer; the same holds for $\mathcal{B}$, which is an increasing transform of $g$. ∎

### v2

@@ 6 Related Work
Hallucination in Language Models. Hallucinations represent a fundamental challenge in modern language models. Comprehensive surveys establish taxonomies that distinguish factuality vs. faithfulness (Huang et al., 2024; Tonmoy et al., 2024). Theoretical work proves hallucinations are mathematically inevitable in computable models under certain resource constraints (Xu et al., 2024; Kalai & Vempala, 2024), with smaller models exhibiting particular severity on factual benchmarks (Lin et al., 2021; Li et al., 2023a). Mechanistic investigations reveal hallucinations arise from internal representation failures (Yu et al., 2024), knowledge awareness limitations (Ferrando et al., 2025), and attention pattern anomalies.
Mitigation has predominantly targeted average performance. Retrieval augmentation (RAG) incorporates external knowledge for factual grounding (Niu et al., 2024). RLHF improves alignment (Bai et al., 2022), while constitutional AI enhances safety. Decoding methods use contrastive decoding (Li et al., 2023b) and classifier-free guidance (Sanchez et al., 2023). Critically, improving $\operatorname*{\mathbb{E}}[\text{error}]$ does not guarantee improvements to $\mathbb{P}(\text{hallucination})$, as tail events depend on variance and correlation structure, not just central tendency.
Second-moment approaches exist but lack theoretical grounding: self-consistency reduces hallucinations through diverse sampling (Wang et al., 2022) without formal tail-probability guarantees, while deep ensembles provide uncertainty estimates (Lakshminarayanan et al., 2017) but not hallucination-specific bounds. We provide the first formal tail bounds connecting neural diversity to hallucination probability as a second-moment problem.
Deep Ensembles, Parallel Architectures & Inference-Time Scaling Deep ensembles provide uncertainty estimates (Lakshminarayanan et al., 2017) with power-law scaling (Lobacheva et al., 2020) for calibration and OOD detection. LLM ensembles benefit from explicit diversity optimization (Tekin et al., 2024), while negative correlation learning demonstrates diversity must be actively encouraged (Liu & Yao, 1999). The “memory split advantage” shows ensembles of smaller models can outperform single large models at fixed parameter budgets. Optimal size theory reveals weighted voting exhibits diminishing returns due to correlation and overfitting (Bonab & Can, 2019), with predictions stabilizing at 5–10 models (Hernández-Lobato et al., 2013). These approaches require multiple independent models, incurring $P\times$ training and inference costs.
Inference-time methods reduce hallucinations through diverse sampling and aggregation. Self-consistency uses majority voting over multiple generations (Wang et al., 2022). Confidence-based weighting uses intelligent aggregation (Taubenfeld et al., 2025), while contrastive decoding contrasts expert and amateur models (Li et al., 2023b). These approaches require multiple forward passes at inference time, whereas our training-time parallelism learns coordinated streams.
Self-ensembled parallel architectures like ParScale (Chen et al., 2025) break the multiplicative memory requirements of classical ensembles by using $P$ perturbed computational pathways within a single model. ParScale achieves $O(\log P)$ general capability gains, modeling parallel streams with correlation $\rho$ in scaling laws $L\propto(N\cdot P^{1/\alpha}\cdot[(P-1)\rho+1]^{-1/\alpha})^{-\alpha}$. This targets mean loss for accuracy improvements, not hallucination probability. We directly build our demonstration on ParScale, extending their theoretical framework and implementation to tail-bound hallucinations.
Theoretical Foundations. Modern portfolio theory (Markowitz, 1952) provides the mathematical foundation for understanding correlation-based risk reduction, with diversification principles (Meucci, 2009) for ensemble variance analysis. Classical ensemble theory reduces mean error $\operatorname*{\mathbb{E}}[\text{loss}]$ via variance decomposition (Dietterich, 2000). PAC-Bayesian bounds connect diversity to minimax-optimal generalization (Ortega et al., 2022) and concentration inequalities showing correlation reduction tightens tail bounds (Alquier, 2024). We link these frameworks to modern neural networks to bound hallucination tail probabilities.
Redundancy Reduction. A rich history of diversification exists in self-supervised learning to avoid training collapse and in PEFT methods for efficient specialization. Self-supervised approaches like Barlow Twins (Zbontar et al., 2021) and VICReg (Bardes et al., 2022) use decorrelation to prevent dimensional collapse (Jing et al., 2022). PEFT methods like LoRA (Hu et al., 2022) and prefix-tuning (Li & Liang, 2021) enable model specialization under limited parameter budgets, with BatchEnsemble and LoRA-Ensemble achieving diversity through parameterization (Wen et al., 2020; Mühlematter et al., 2025). We adapt these methods for second-moment reliability guarantees.
@@ 7 Discussion
At a time when the reliability of language models is becoming the critical barrier to real-world deployment, we (i) provide the first formal framework to tail-bound hallucinations in ensembled language models, demonstrating that neural diversity plays a critical role in reducing hallucinations; and, (ii) using this technique, achieve up to 25.6% (and 14.6% on average) reduction in hallucination rates at fixed parameter and data budgets at +0.004% pretraining cost. Neural diversity enables reliability gains without massive compute scaling.
By reframing hallucinations as a second-moment problem — controlled through variance and correlation rather than mean optimization — we open an under-explored research direction orthogonal to existing approaches. While RLHF and RAG target first-moment improvements (average performance), neural diversity targets tail probability through explicit decorrelation. This bridges portfolio theory to neural reliability, a connection previously unexplored. The gap between extensive first-moment research and nascent second-moment approaches (self-consistency, our work) suggests substantial opportunity for reliability-focused methods grounded in tail-probability theory.
Our small-scale demonstration and mechanistic analysis validates the theoretical framework; scaling to production models is straightforward given that continued training requires only +0.004% additional overhead and $P=4$ captures 94.3% of oracle performance. The task-dependent optimal $P_{\star}$ in Table 1 reveals intriguing structure, suggesting deeper connections between task complexity, knowledge recall vs. precision and neural diversity worthy of theoretical characterization.
Our work opens two immediate research directions: (i) Theoretical: characterizing optimal $P_{\star}$ as a function of task properties — our U-shape theorem (Theorem 2) suggests information-theoretic approaches. (ii) Practical: combining neural diversity (this work) with inference-time scaling (Snell et al., 2024) for multiplicative reliability gains. Second-moment reliability is an essential frontier as LLMs become critical infrastructure in high-stakes domains.
@@ Acknowledgments
We thank the research community at South Park Commons for valuable discussions and feedback throughout this project. We are grateful to Augustine Mavor-Parker, Bhav Ashok, Daniel Morillo, David Sontag, Iman Modarressi, Jaclyn Lunger, Javier Ferrando, John Bohannon, Nelson Ray, Patrick Bozeman and Suhith Rajesh for their insightful comments on earlier drafts. This work was supported by South Park Commons and Modal, who generously provided compute resources for our experiments.
@@ A.1.1 Proof of Lemma 1
We proceed in two steps. First, we show that the shared linear readout $A$ can at most distort cosines between representations by a factor of $\kappa^{2}$. Second, we convert bounded cosine alignment between the error vectors $\xi_{i}$ into a bound on average noise correlation.
Step 1: local linear readout and cosine distortion.
Let $\sigma_{\min}$ and $\sigma_{\max}$ denote the minimal and maximal singular values of $A$ and $\kappa=\sigma_{\max}/\sigma_{\min}$ its condition number. For any nonzero $x,y\in\mathbb{R}^{d}$, set $u=Ax$ and $v=Ay$. Then
$\cos\angle(u,v)\;=\;\frac{\langle u,v\rangle}{\|u\|_{2}\,\|v\|_{2}}\;=\;\frac{x^{\top}A^{\top}Ay}{\|Ax\|_{2}\,\|Ay\|_{2}}.$
By the spectral bounds on $A^{\top}A$, we have
$|x^{\top}A^{\top}Ay|\;\leq\;\|A^{\top}A\|_{\mathrm{op}}\,|x^{\top}y|\;=\;\sigma_{\max}^{2}\,|x^{\top}y|$
and
$\|Ax\|_{2}\,\|Ay\|_{2}\;\geq\;\sigma_{\min}^{2}\,\|x\|_{2}\,\|y\|_{2}.$
Combining,
$|\cos\angle(u,v)|\;\leq\;\frac{\sigma_{\max}^{2}}{\sigma_{\min}^{2}}\frac{|x^{\top}y|}{\|x\|_{2}\,\|y\|_{2}}\;=\;\kappa^{2}\,|\cos\angle(x,y)|.$
Squaring both sides yields
$\cos^{2}\angle(u,v)\;\leq\;\kappa^{4}\,\cos^{2}\angle(x,y).$ (10)
Apply this with $x=\tilde{z}_{i}(X)$ and $y=\tilde{z}_{j}(X)$. Under the local linear readout assumption we have $\xi_{i}(X)=A\tilde{z}_{i}(X)$ and $\xi_{j}(X)=A\tilde{z}_{j}(X)$, so $u=\xi_{i}(X)$ and $v=\xi_{j}(X)$. Thus, for every pair $(i,j)$ and every input $X$ with nonzero norms,
$\cos^{2}\angle\big(\xi_{i}(X),\xi_{j}(X)\big)\;\leq\;\kappa^{4}\,\cos^{2}\angle\big(\tilde{z}_{i}(X),\tilde{z}_{j}(X)\big).$
Taking expectations over $X$ gives
$\mathcal{D}_{\xi,ij}^{2}\;\triangleq\;\operatorname*{\mathbb{E}}_{X}\big[\cos^{2}\angle(\xi_{i}(X),\xi_{j}(X))\big]\;\leq\;\kappa^{4}\,\operatorname*{\mathbb{E}}_{X}\big[\cos^{2}\angle(\tilde{z}_{i}(X),\tilde{z}_{j}(X))\big]\;\triangleq\;\kappa^{4}\,\mathcal{D}_{ij}^{2},$
where $\mathcal{D}_{ij}^{2}$ denotes the pairwise cosine diversity in representation space. Averaging over pairs and taking square roots yields
$\mathcal{D}_{\xi}\;\triangleq\;\sqrt{\operatorname*{\mathbb{E}}_{i<j}\mathcal{D}_{\xi,ij}^{2}}\;\leq\;\kappa^{2}\,\sqrt{\operatorname*{\mathbb{E}}_{i<j}\mathcal{D}_{ij}^{2}}\;=\;\kappa^{2}\,\mathcal{D}.$ (11)
Step 2: from cosine alignment to average correlation.
We now bound the average correlation $\bar{\rho}$ in terms of $\mathcal{D}_{\xi}$. Fix a pair $(i,j)$ and write
$X\;\triangleq\;\langle\xi_{i},\xi_{j}\rangle,\qquad B\;\triangleq\;\|\xi_{i}\|_{2}\,\|\xi_{j}\|_{2}.$
Whenever $B>0$,
$\cos\angle(\xi_{i},\xi_{j})\;=\;\frac{X}{B},\qquad\mathcal{D}_{\xi,ij}^{2}\;=\;\operatorname*{\mathbb{E}}\!\left[\left(\frac{X}{B}\right)^{2}\right].$
Using Cauchy–Schwarz with $U=X/B$ and $V=B$, we obtain
$\Sigma_{ij}^{2}=\big(\operatorname*{\mathbb{E}}[X]\big)^{2}=\big(\operatorname*{\mathbb{E}}[UV]\big)^{2}\;\leq\;\operatorname*{\mathbb{E}}[U^{2}]\,\operatorname*{\mathbb{E}}[V^{2}]=\mathcal{D}_{\xi,ij}^{2}\,\operatorname*{\mathbb{E}}\big[\|\xi_{i}\|_{2}^{2}\,\|\xi_{j}\|_{2}^{2}\big].$
Apply Cauchy–Schwarz again to the norms and use the kurtosis bound:
$\operatorname*{\mathbb{E}}\big[\|\xi_{i}\|_{2}^{2}\,\|\xi_{j}\|_{2}^{2}\big]\;\leq\;\sqrt{\operatorname*{\mathbb{E}}[\|\xi_{i}\|_{2}^{4}]\,\operatorname*{\mathbb{E}}[\|\xi_{j}\|_{2}^{4}]}\;\leq\;\sqrt{C_{4}\sigma_{i}^{4}\,C_{4}\sigma_{j}^{4}}=C_{4}\,\sigma_{i}^{2}\sigma_{j}^{2}.$
Combining,
$\Sigma_{ij}^{2}\;\leq\;C_{4}\,\mathcal{D}_{\xi,ij}^{2}\,\sigma_{i}^{2}\sigma_{j}^{2},\qquad\rho_{ij}^{2}=\frac{\Sigma_{ij}^{2}}{\sigma_{i}^{2}\sigma_{j}^{2}}\;\leq\;C_{4}\,\mathcal{D}_{\xi,ij}^{2},$
so
$|\rho_{ij}|\;\leq\;\sqrt{C_{4}}\,\mathcal{D}_{\xi,ij}.$ (12)
Finally, average over pairs and apply Cauchy–Schwarz in the index space:
$|\bar{\rho}|=\left|\operatorname*{\mathbb{E}}_{i<j}[\rho_{ij}]\right|\;\leq\;\operatorname*{\mathbb{E}}_{i<j}[|\rho_{ij}|]\;\leq\;\sqrt{C_{4}}\,\operatorname*{\mathbb{E}}_{i<j}[\mathcal{D}_{\xi,ij}]\;\leq\;\sqrt{C_{4}}\,\sqrt{\operatorname*{\mathbb{E}}_{i<j}\mathcal{D}_{\xi,ij}^{2}}=\sqrt{C_{4}}\,\mathcal{D}_{\xi}.$
Plugging equation 11 into this inequality gives
$|\bar{\rho}|\;\leq\;\sqrt{C_{4}}\,\mathcal{D}_{\xi}\;\leq\;\sqrt{C_{4}}\,\kappa^{2}\,\mathcal{D}\;=\;C_{\ast}\,\mathcal{D},$
with $C_{\ast}=\sqrt{C_{4}}\,\kappa^{2}$, as claimed. ∎
A.1.2 Proof of Theorem 1
@@
We work under the signal–noise model from the preliminaries. By Markowitz 1952,
$\mathrm{Var}\left(\frac{1}{P}\sum_{i=0}^{P-1}X_{i}\right)=\bar{\sigma}^{2}\left(\frac{1-\rho}{P}+\rho\right)$
where $\bar{\sigma}^{2}=\frac{1}{P}\sum_{i}\mathrm{Var}(X_{i})$ is the average variance and $\rho=\operatorname*{\mathbb{E}}_{i\neq j}[\mathrm{Corr}(X_{i},X_{j})]$ is the average pairwise correlation.
From Lemma 1, we have
$|\bar{\rho}|\;\leq\;C_{\ast}\,\mathcal{D}.$
Substituting this into the variance expression yields
$\mathrm{Var}(E_{w})$ $\leq\bar{\sigma}^{2}\left(\frac{1-C_{\ast}\,\mathcal{D}}{P}+C_{\ast}\,\mathcal{D}\right),$
which is exactly the claimed variance bound in Theorem 1.
By construction, the hallucination event is
$\mathrm{H}_{\delta}\;\triangleq\;\{E_{w}\geq\delta\},\qquad\delta>0,$
and we have already noted that $\operatorname*{\mathbb{E}}[E_{w}]=0$. Applying the one-sided Chebyshev inequality from the preliminaries to the random variable $E_{w}$ with mean 0 and variance $v=\mathrm{Var}(E_{w})$ gives
$\mathbb{P}(\mathrm{H}_{\delta})\;=\;\mathbb{P}(E_{w}\geq\delta)\;\leq\;\frac{v}{v+\delta^{2}}\;=\;\frac{\mathrm{Var}(E_{w})}{\mathrm{Var}(E_{w})+\delta^{2}}.$
Substituting $\mathrm{Var}(E_{w})$ by its upper bound yields
$\mathbb{P}(\mathrm{H}_{\delta})\;\leq\;\frac{\bar{\sigma}^{2}\left(\frac{1-C_{\ast}\,\mathcal{D}}{P}+C_{\ast}\,\mathcal{D}\right)}{\bar{\sigma}^{2}\left(\frac{1-C_{\ast}\,\mathcal{D}}{P}+C_{\ast}\,\mathcal{D}\right)+\delta^{2}}.$
Dividing both numerator and denominator by $\bar{\sigma}^{2}$ matches the bound stated in Theorem 1. ∎
@@ A.1.3 Proof of Theorem 2
Extend $P$ to a real variable with domain $P\geq 1$; the claim for integer $P$ follows by restriction.
Under uniform weights $w_{i}=1/P$, the ensemble error variance can be written as
$v(P)\triangleq\mathrm{Var}(E_{w})=\bar{\sigma}^{2}\left(\frac{1-\bar{\rho}(P)}{P}+\bar{\rho}(P)\right),$
with $\bar{\sigma}^{2}>0$ and
$\bar{\rho}(P)=\rho_{0}+\beta(P-1)^{\gamma},\qquad\rho_{0}\in[0,1),\;\beta>0,\;\gamma>0.$
The bound from the main text is
$\mathbb{P}(\mathrm{H}_{\delta})\leq B(P)\triangleq\frac{v(P)}{v(P)+\delta^{2}},\qquad\delta>0.$
Step 1: Reduction to $v(P)$.
Define $\phi(x)\triangleq x/(x+\delta^{2})$ for $x\geq 0$. Then
$\phi^{\prime}(x)=\frac{\delta^{2}}{(x+\delta^{2})^{2}}>0,$
so $\phi$ is strictly increasing. Hence $B(P)=\phi\big(v(P)\big)$ has the same extrema and monotonicity as $v(P)$. Since $\bar{\sigma}^{2}>0$, it suffices to analyze
$f(P)\triangleq\frac{v(P)}{\bar{\sigma}^{2}}=\frac{1-\bar{\rho}(P)}{P}+\bar{\rho}(P).$
Step 2: First derivative and unique critical point.
For $P>1$,
$\bar{\rho}^{\prime}(P)=\beta\gamma(P-1)^{\gamma-1}.$
A direct calculation gives
$f^{\prime}(P)$ $=\frac{\mathrm{d}}{\mathrm{d}P}\Big(\tfrac{1-\bar{\rho}(P)}{P}+\bar{\rho}(P)\Big)$
$=\frac{\beta(P-1)^{\gamma}(P\gamma+1)+(\rho_{0}-1)}{P^{2}}\triangleq\frac{N(P)}{P^{2}}.$
We study $N(P)$.
At $P=1$ we have
$N(1)=\beta\cdot 0^{\gamma}(\gamma+1)+(\rho_{0}-1)=\rho_{0}-1<0.$
Differentiating $N$ for $P>1$ yields
$N^{\prime}(P)=\beta\gamma(\gamma+1)\cdot P\cdot(P-1)^{\gamma-1}.$
All factors on the right are strictly positive for $P>1$, so $N^{\prime}(P)>0$ on $(1,\infty)$ and $N$ is strictly increasing. Moreover,
$(P-1)^{\gamma}(P\gamma+1)\sim\gamma P^{\gamma+1}\xrightarrow[P\to\infty$$\infty$,
so $N(P)\to+\infty$ as $P\to\infty$. By continuity and strict monotonicity, there exists a unique $P_{\star}>1$ such that $N(P_{\star})=0$.
Because $P^{2}>0$ for all $P\geq 1$, the sign of $f^{\prime}(P)$ matches that of $N(P)$:
$f^{\prime}(P)\begin{cases}<0,&1<P<P_{\star},\\[2.0pt] =0,&P=P_{\star},\\[2.0pt] >0,&P>P_{\star}.\end{cases}$
Thus $f$ (and hence $v$) is strictly decreasing on $(1,P_{\star})$ and strictly increasing on $(P_{\star},\infty)$; $P_{\star}$ is the unique global minimizer.
Step 3: U-shape of the hallucination bound.
Since $B(P)=\phi\big(v(P)\big)$ and $\phi$ is strictly increasing,
$B^{\prime}(P)=\phi^{\prime}(v(P))\cdot v^{\prime}(P),\qquad\phi^{\prime}(v(P))>0,$
so $B$ inherits the same monotonicity: it is strictly decreasing on $(1,P_{\star})$, strictly increasing on $(P_{\star},\infty)$, and
$B(P_{\star})=\min_{P\geq 1}B(P).$
Therefore the upper bound on $\mathbb{P}(\mathrm{H}_{\delta})$ is U-shaped in $P$ with a unique global minimum at $P_{\star}$, determined by the parameters $(\rho_{0},\beta,\gamma)$ governing $\bar{\rho}(P)$. ∎